Shift report 20000826 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Aug 26 2000 - 07:40:40 EDT

  • Next message: http account: "Shift report 20000826 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Aug 26 07:40:40 2000, Shift: 00:00-08:00
    Supervisor: ZM, with: RO
    Log book#5 pages 18 - 19
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam stored and cogged at 1:35. Smooth runing up to 6:15
       when beam has been dumped. Beam intesity lower than 
       during former shifts: triger # 1 = (10 - 20)/sec.
    
    2. Two runs recorded (#2443 with 341877 events #2444 with 
       172009 events).
       MRD at 45 deg., D5(700A, polarity A)
       FS  at  5 deg., D1(625A, polarity B) D2(492A, polarity B)
    
    3. No problems 
    
    
      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Aug 26 2000 - 07:44:03 EDT